• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name ulmer

    Log entry time 16:05:45 on June13,2002

    Entry number 86001

    keyword=Shift summary

    Start shift with continuation of Q2D_b50 point.

    10:15 ROC14 and ROC 1 crashed. Declare the Q2D_b50 point done:
    Got 300 of the 330 wanted coincidence events.
    10:30 DAQ back up and running: do Q2D_b50 efficiency run.
    10:35 MCC takes beam away for injector studies. They estimate minimum
    down time of 1 hour. Set up for Q2D_bcN in the meantime.
    13:25 Beam back (low current). A perfect opportunity to do Arc energy
    measurement since only 5 uA is needed and that's what we now have.
    13:30 While waiting for Arun, continue with production data: Q2D_bcN
    at 50 uA. More DAQ problems, so this point takes a bit longer than
    expected.
    14:15 End Q2D_bcN and set up for Arc energy measurement: 5 uA unrastered
    on optics target. OTR near target shows multiple peaks in Y for unrastered
    beam (also saw signal outside of normal raster pattern previously - just
    before MCC took beam away at 10:35 and until 14:15). Arun is talking
    with MCC now to troubleshoot.
    14:30 Now OTR looks fine (unrastered shows 1 narrow peak).
    14:31 Arc energy measurement fails due to harp scanner problem. This will
    need to be dealt with during next development day. May entail bringing
    that section of beamline up to air and replacing module (few hours).
    See Halog 85999.
    14:40 Harp scan (performed by Arun). See Halog 86000. OTR widths are
    twice as large as Harp, roughly, in both X and Y. MCC is looking into
    it.
    15:00 Bodo starts ep energy measurement.
    16:00 Bodo finishes ep measurement. I hear thunder and the beam goes away.
    I wonder if there is a connection.

    We are currently set up for the Q2D_c20 kinematics, though no data has been
    taken yet on this point.